# Quantum simulation of one-dimensional fermionic systems with Ising Hamiltonians

### v1

@@ VIII Conclusion and outlook
A perturbative analysis of the error due to the approximation made in our method showed that the main source of error are perturbations of the energy eigenvalues of the $M$-particle domain wall Hamiltonian $H_{DW}^{(M)}$ leading to phase deviations. These perturbations are caused by interactions with states in subspaces of distinct $M$. The analysis shows that the error can be made arbitrarily small by increasing the strength of the domain wall coupling $|J|$, or alternatively by re-scaling the model parameters at moderate a time-overhead. The re-scaling of the model parameters leads to a time complexity of our continuous time simulation method that scales with the square root of the inverse error $1/\varepsilon$. This is more efficient than first-order product formulas for gate-based quantum simulators, where the error scales with $1/\varepsilon$. Therefore, our approach demonstrates an advantage of programmable Ising Hamiltonians as analog quantum simulators of fermionic systems over their gate-based counterparts in terms of worst-case error.
In order to gauge the viability of an implementation of our approach in real quantum hardware, we investigate the impact of decoherence on the simulation fidelity and find that for realistic estimates of the single-qubit relaxation and dephasing error rates, dynamic properties of the Aubry-Andre model can be observed for short time evolutions. This makes our approach a reasonable benchmarking task to compare analog NISQ-devices. Our method allows to reproduce various recent results for the aforementioned models using only a simple Ising Hamiltonian. This moves experimental verification of those results on programmable quantum hardware within reach. This will be examined in more detail in the future.
Our work suggests various future lines of research, most notably towards the experimental realization of the systems we have discussed here. Beyond that, we believe it will be insightful to further investigate the possibilities of Floquet engineering in order to realize longer ranging interactions, as well as higher dimensional effective models. Furthermore, while we have investigated both the SSH model, as well as an interacting XXZ chain, our approach also allows to combine the SSH model with nearest-neighbor interactions. This model has been shown to exhibit a rich phase diagram [59]. Additionally, the ability of simple Ising Hamiltonians to simulate spin chains makes them a viable candidate as a component in larger quantum information processing architectures. Thus, in the future we intend to examine the quantum state transfer properties of the approach more closely.
@@ Appendix A Symmetries and consistent simulations
As discussed above, the Hamiltonian $H_{DW}$, as well as any other operator $O$ that is consistent with the domain wall encoding, commutes with $\Pi$. Therefore, the Hilbert space is now divided into two subspaces of equal dimension, spanned by eigenstates of $\Pi$ with eigenvalue ±$\pm 1$ respectively. Consistent operators act the same upon each subspace. In fact, $H_{DW}$ is equivalent to $H_{\text{Fermi}}$ in both subspaces, so one way to mitigate the non-uniqueness of the mapping from fermion states to domain wall states could be to restrict to either of the ±$\pm 1$-eigenspaces of $\Pi$ and the fermionic state could be mapped to domain wall states according to
@@ Appendix B Mapping operators to domain wall encoding
The minimum and maximum in the indices generalizes to cases where $i>j$. As a convention, if the index $n$ does not run over positive numbers, $X_{ij}=\mathbb{I}$, which is particularly the case if $i=j$. Whether a domain wall is created or destroyed depends on there already being domain walls on the respective site. Consequently, the operator $X_{ij}$ acting on the Ising chain can be identifies with the spin chain operator
@@
Using the mapping Eq. (100) in conjunction with the Jordan-Wigner transformation, any Hermitian operator acting on spins or fermions can be expressed as an operator acting on domain walls, with the restriction that the decomposition Eq. (95) only contains operator strings of even length. While this seems like a substantive constraint at first glance, there are two things to consider for context. First, this is in line with the parity superselection rule for fermionic systems [60] and, second, our approach effectively models a class of fermionic Hamiltonian that preserves the particle number. Naturally, this preserves also the particle number parity and expectation values implicitly project any operator in the respective parity subspace. Therefore, any physical observable can be mapped to the domain wall picture.

### v2

@@ VIII Conclusion and outlook
A perturbative analysis of the error due to the approximation made in our method showed that the main source of error are perturbations of the energy eigenvalues of the $M$-particle domain wall Hamiltonian $H_{DW}^{(M)}$ leading to phase deviations. These perturbations are caused by interactions with states in subspaces of distinct $M$. The analysis shows that the error can be made arbitrarily small by increasing the strength of the domain wall coupling $|J|$, or alternatively by re-scaling the model parameters at moderate a time-overhead. The re-scaling of the model parameters leads to a time complexity of our continuous time simulation method that scales the same in terms of time and error as first-order product formulas for gate-based quantum simulators.
In order to gauge the viability of an implementation of our approach in real quantum hardware, we investigate the impact of decoherence on the simulation fidelity and find that for realistic estimates of the single-qubit relaxation and dephasing error rates, dynamic properties of the Aubry-Andre model can be observed for short time evolutions. This makes our approach a reasonable benchmarking task to compare analog NISQ-devices. Our method allows to reproduce various recent results for the aforementioned models using only a simple Ising Hamiltonian. This moves experimental verification of those results on programmable quantum hardware within reach and will be examined in more detail in the future.
Our work suggests various future lines of research, most notably towards the experimental realization of the systems we have discussed here, e.g. using superconducting qubits or Rydberg atoms. Beyond that, we believe it will be insightful to further investigate the possibilities of Floquet engineering in order to realize longer ranging interactions, as well as higher dimensional effective models. Furthermore, while we have investigated both the SSH model, as well as an interacting XXZ chain, our approach also allows to combine the SSH model with nearest-neighbor interactions. This model has been shown to exhibit a rich phase diagram [59]. Additionally, the ability of simple Ising Hamiltonians to simulate spin chains makes them a viable candidate as a component in larger quantum information processing architectures. Thus, in the future we intend to examine the quantum state transfer properties of the approach more closely.
@@ Appendix A Symmetries and consistent simulations
The Hamiltonian $H_{DW}$, as well as any other operator $O$ that is consistent with the domain wall encoding, commutes with $\Pi$. Therefore, the Hilbert space is now divided into two subspaces of equal dimension, spanned by eigenstates of $\Pi$ with eigenvalue ±$\pm 1$ respectively. Consistent operators act the same upon each subspace. In fact, $H_{DW}$ is equivalent to $H_{\text{Fermi}}$ in both subspaces, so one way to mitigate the non-uniqueness of the mapping from fermion states to domain wall states could be to restrict to either of the ±$\pm 1$-eigenspaces of $\Pi$ and the fermionic state could be mapped to domain wall states according to
@@ Appendix B Mapping operators to domain wall encoding
The minimum and maximum in the indices generalize to cases where $i>j$. As a convention, if the index $n$ does not run over positive numbers, $X_{ij}=\mathbb{I}$, which is particularly the case if $i=j$. Whether a domain wall is created or destroyed depends on there already being domain walls on the respective site. Consequently, the operator $X_{ij}$ acting on the Ising chain can be identified with the spin chain operator
@@
Using the mapping Eq. (100) in conjunction with the Jordan-Wigner transformation, any Hermitian operator acting on spins or fermions can be expressed as an operator acting on domain walls, with the restriction that the decomposition Eq. (95) only contains operator strings of even length. While this seems like a substantive constraint at first glance, there are two things to consider for context. First, this is in line with the parity superselection rule for fermionic systems [60] and, second, our approach effectively models a class of fermionic Hamiltonian that preserves the particle number. Naturally, this preserves also the particle number parity and expectation values implicitly project any operator in the respective parity subspace. Therefore, any relevant physical observable can be mapped to the domain wall picture.
@@
Appendix C Relationship between $\varepsilon_{\|\cdot\|}$ and $\varepsilon$
As discussed in the main text, we computed the error in terms of the infidelity of the simulation, which we denote as $\varepsilon$. The error used in the gate complexity of discretized simulations of continuous time evolutions is formulated in terms of the spectral norm of the distance between the respective unitaries [38], which we denote as $\varepsilon_{\|\cdot\|}$. In order to compare the error scaling between Trotterization and our analog approach, we need to compare these two error metrics.
In the main text, we denote the target unitary as $U_{T}$, while its approximation is denoted with $\tilde{U}_{T}$, which we can decompose into the target unitary and an error as
$\tilde{U}_{T}=U_{\varepsilon}U_{T}$ (102)
By choosing an appropriate basis, we can compare the error measured in terms of the fidelity and the distance without loss of generality by considering the identity operator $\mathbb{I}$ and the error unitary $U_{\varepsilon}=\sum_{n}e^{-i\phi_{n}}|\phi_{n}\rangle\langle\phi_{n}|$. Furthermore, since we are interested in the asymptotic behaviour for small errors, we will assume the arguments of the eigenvalues of $U_{\varepsilon}$ to be small, i.e. $|\phi_{n}|\ll 1$. We first calculate the spectral norm of the distance
$\varepsilon_{\|\cdot\|}$ $=\|\mathbb{I}-U_{\varepsilon}\|$ (103)
$=\sqrt{\lambda_{\max}((\mathbb{I}-U_{\varepsilon})^{\dagger}(\mathbb{I}-U_{\varepsilon}))}$
$=\sqrt{\lambda_{\max}(2\mathbb{I}-U^{\dagger}_{\varepsilon}-U_{\varepsilon})}$
$=\sqrt{\max_{n}\{2-2\cos\phi_{n}\}}$
$\approx|\phi_{\max}|$
where $\lambda_{\max}(O)$ denotes the maximum eigenvalue of the operator $O$ and $\phi_{\max}:=\max_{n}\phi_{n}$. Equivalently, we can compute the fidelity
$\mathcal{F}$ $=|\langle\Psi_{0}|U_{T}^{\dagger}\tilde{U}_{T}|\Psi_{0}\rangle|^{2}=|\langle\Psi_{T}|U_{\varepsilon}|\Psi_{T}\rangle|^{2}$ (104)
$=\left|\sum_{n}|\tilde{a}_{n}|^{2}e^{-i\phi_{n}}\right|^{2}$
$=\sum_{m,n}|\tilde{a}_{n}|^{2}|\tilde{a}_{m}|^{2}e^{i(\phi_{m}-\phi_{n})}$
$=\sum_{n}|\tilde{a}_{n}|^{4}+2\sum_{m>n}|\tilde{a}_{m}|^{2}|\tilde{a}_{n}|^{2}\cos(\phi_{m}-\phi_{n})$
where we write $U_{T}|\Psi_{0}\rangle=|\Psi_{T}\rangle=\sum_{n}\tilde{a}_{n}|\phi_{n}\rangle$ in the eigenbasis of the error $U_{\varepsilon}$. As before, we consider small errors and the above expression can be approximated as
$\mathcal{F}$ $\approx\sum_{n}|\tilde{a}_{n}|^{4}$ (105)
$+2\sum_{m>n}|\tilde{a}_{m}|^{2}|\tilde{a}_{n}|^{2}\left(1-\frac{1}{2}(\phi_{m}-\phi_{n})^{2}\right)$
$=1-\sum_{n,m>n}|\tilde{a}_{m}|^{2}|\tilde{a}_{n}|^{2}(\phi_{m}-\phi_{n})^{2}$
and subsequently lower bounded by
$\mathcal{F}$ $\geq 1-4\phi_{\max}^{2}\sum_{n,m>n}|\tilde{a}_{m}|^{2}|\tilde{a}_{n}|^{2}$ (106)
$=1-2\phi_{\max}^{2}\left(1-\sum_{n}|\tilde{a}_{n}|^{4}\right)\ .$
The lower bound on the fidelity corresponds to the maximum of the eigenvalues considered in the operator norm in Eq. (103).
From Eq. (103), Eq. (106) and with the definition of the infidelity as $\epsilon=1-\mathcal{F}$, we find that $\varepsilon\leq O(\phi_{\max}^{2})=O\left(\varepsilon_{\|\cdot\|}^{2}\right)$. As discussed in the main text, this shows that our continuous time quantum simulation algorithm has the same error scaling as first-order Trotterization.
[…]
$=\frac{1}{d^{2}}\left(d+\sum_{m>n}2\cos(\phi_{n}-\phi_{m})\right)\ .$
As before, this can be bounded using the maximum $\phi_{\max}$ and subsequently expanded for small arguments as
$\mathcal{F}_{op}$ $\geq\frac{1}{d^{2}}\left(d+(d^{2}-d)\cos 2\phi_{\max}\right)$ (108)
$\approx\frac{1}{d}+\left(1-\frac{1}{d}\right)(1-2\phi_{\max}^{2})$
$=1-2\phi_{\max}^{2}\left(1-\frac{1}{d}\right)\ .$
where in the approximation we use the Taylor expansion of the cosine. Hence, for small errors, the infidelity using this alternative definition is also $\varepsilon\leq O(\phi_{\max}^{2})$.
The sum $\sum_{n}|\tilde{a}_{n}|^{4}$ can be bounded according to
$\frac{1}{d}\leq\sum_{n}|\tilde{a}_{n}|^{4}\leq 1\ ,$ (109)
which can easily be shown using the Cauchy-Schwarz inequality. Hence, the fidelity of a specific initial state is lower bounded by the operator fidelity. Furthermore, the sum $\sum_{n}|\tilde{a}_{n}|^{4}$ is related to the localization of the evolved state in the eigenbasis of the error via the participation ratio. In the main text we have argued that the leading error is caused by perturbations on the energy eigenvalues. This effectively translates to an error $U_{\varepsilon}$ that is diagonal in the eigenbasis of the target Hamiltonian. Thus, the calculation here provides an alternative argument why extended states in the eigenbasis of the time-evolution operator are more affected by errors than localized states.